A fastfood chain's softdrink machine is calibrated so that it disperses an average of 225 milliliters per cup. The amount of drink is normally distributed with a standard deviation equal to 15 milliliters.


1
Expert's answer
2021-11-03T17:56:53-0400

Question is incomplete. We can add here:

A fast-food chain's soft-drink machine is calibrated so that it disperses an average of 225 milliliters per cup. The amount of drink is normally distributed with a standard deviation equal to 15 milliliters. How many cups will probably overflow if 255- milliliter cups are used for the next 1000 drinks?

Solution:

Let X represent the amount of drink distributed.

"\mu = 225 \\

\sigma=15"

The number of cups that will probably overflow if 230-milliliter cups are used for the next 1000 drinks

"P(X>230) = 1 -P(X<230) \\

= 1 -P(Z< \frac{255-225}{15}) \\
= 1 -P(Z< \frac{30}{15}) \\

= 1 -P(Z<2) \\

= 1 -0.9772 \\

= 0.0228 \\

E(X) = n \times P \\

=1000 \times 0.0228 \\

= 22.8 ≈ 23"

Hence, 23 cups will overflow.